The value of cos a/sin(90+a)  +sin a/(sin (180+a)  +cos (90+a)/tan a

Grade:10

1 Answers

Arun
6 years ago
As we know
Sin(90+A) = cos A
Sin(180+A) = - sinA
Cos(90+A) = -sin A
Put thus conversion and find value